Given the two points(4,6) and (8,18):
1. Calculate the slpoe.
2. Write the point-slope form of the line through the two points.
3. Use the answer to #9 to write the equation in slope-intercept form.

Answer by Mathtut(3670)

slope is change in y over change in x
:
(18-6)/(8-4)=12/4=3
:
so slope is 3
:
use point slope formula....using slope 3 and either given point...I am using (4,6)
:
y-6=3(x-4)
:
y-6=3x-12
:
y=3x-6 in slope intercept form